Sat Tranh took a short hike with his friends up Mt. Wachusett. They hiked uphill at a steady pace of 1.5 mph, and downhill at a rate of 4mph. If the time to climb the mt. took an hour more than the time to hike down, how long was the entire hike?
I tried doing d=rt but there is no distance given here. I also tried 4x= 1.5x-1.5 but it yields the incorrect answer. If someone could help me understand how to do this I would greatly appreciate it!

Found 2 solutions by mananth, scott8148:
Answer by mananth(16946)

Let the time to hike down be x
The time to hike up was x+1 hours
distance down = 4x
distance up = 1.5(x+1)
Both distances are same
1.5x +1.5 =4x
-2.5x = -1.5
x=-1.5/-2.5
x=0.6
x+1=1.6 hours
Total time = 2.2 hours or 2 hours 12 minutes

Answer by scott8148(6628)

if x equals time down, then x+1 equals time up ___ total time is 2x+1

distance is the same, up and down

4 (x) = 1.5 (x + 1) ___ 4x = 1.5x + 1.5

not sure why your equation has the constant term negative...